OpenStudy (anonymous):

i need you guys to check these 5 answers and tell me which are wrong ? I THINK ITS A What is the value of expression? 5^−8 • 5^−4 A. 5^32 B. 5^4 C. 5^−4 D. 5^−12

OpenStudy (dan815):

\[5^{−8} • 5^{−4} = 5^{(-8)+(-4)}=5^{-12}\]

OpenStudy (dan815):

\[A^n * A^m = A^{n+m}\] exponent rules

OpenStudy (anonymous):

c?

OpenStudy (dan815):

Here is an example to give you an idea of this exponent rule \[3^3\] means 3 is multiplying itself 3 times 3*3*3 \[3^3 * 3^2\] means 3 is multiplying itself 3 times and another 3 is multiplying itself 2 times \[(3*3*3) (3*3)\] but this is the same as \[3*3*3*3*3\] so its just 5 3s, in the end its the addition of the exponents \[3^5\]

OpenStudy (anonymous):

ik that

OpenStudy (dan815):

its D

OpenStudy (anonymous):

@dan815

OpenStudy (cookie_2046):

https://mathway.com/ You This To Help You

OpenStudy (dan815):

\[(A^n)^m=A^{nm}\\ A^n *A^m=A^{n+m}\\ \frac{A^n}{A^m}=A^{n-m}\\ A^{-n}=\frac{1}{A^n}\\ \large {A^\frac {n}{m}= ~~~^m\sqrt{A^n}}~~~~m^{th} ~~square-root~~~and~~n^{th}~~power\]

OpenStudy (dan815):

use these rules to start simplifying the expression, what do you get?

OpenStudy (dan815):